OpenStudy (anonymous):

How can sqr(1/6) - 1 be simplified further

OpenStudy (anonymous):

sqr?

OpenStudy (tkhunny):

It can't be simplified, but some manipulation is available. \(\sqrt{1/6} - 1 = \dfrac{1}{\sqrt{6}} - 1 = \dfrac{\sqrt{6}}{6} - 1 = \dfrac{\sqrt{6} - 6}{6}\)